Circular No. 5862

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA SAGITTARII 1993 Independent photographic discoveries of an apparent nova have been

reported by Matsuo Sugano, Minami-Oda, Hyogo, Japan (via S. Nakano, Sumoto), and by William Liller, Viña del Mar, Chile. The object is located at α = 18^{h}09^{m}.7, δ = −29^{o}30′ (equinox 1950.0), situated approximately midway between, and slightly to the north of, two stars of mag 10. Magnitude estimates: Aug. 27.02 UT, [11.5 (Liller; Tech Pan film, orange filter); 14.508, 9.0 (Sugano; Tri-X film); 16.043, 7.9 (Liller).

SUPERNOVA 1987A IN THE LARGE MAGELLANIC CLOUD J. Xu and A. P. S. Crotts, Columbia University; and W. E. Kunkel,

Carnegie Observatories, report: “We have discovered two new (and possibly related) light echo structures from SN 1987A. These are due to material more than twice as distant from the supernova as any previously reported (IAUC 4561, 4733, 4739, 4791), at 3360 and 3470 lyr in the ‘supernova foreground’, corresponding to radii of ∼ 240′′ on 1993 May 1. These first appeared in data from 1991 Sept. 18, at 110^{o} < p.a. < 140^{o} and 3^{o} < p.a. < 13^{o}, respectively. Since then, the second feature has brightened, while the first brightened then later dimmed. Their maximum surface brightness is ∼ 9.2 × 10^{−19} erg s^{−1} cm^{−2} A^{−1} arcsec^{−2} (at 612 nm), and they can be easily detected only once background stars and nebulosity have been subtracted from the image. We also note that the outer prominent echo ring has become very complex, particularly over its western half, with components ranging from ∼ 880 to 2060 lyr distance in the supernova foreground. Further observations of these features are encouraged and would help reveal their continued evolution and therefore their three-dimensional structure.”

SUPERNOVA 1993J IN NGC 3031 CCD photometry by X. Zhou, Y. Li, and H. J. Yan, obtained at the

Xing-Long station of Beijing Astronomical Observatory: Sept. 4.41 UT, B = 14.90, V = 14.32, R = 13.46, I = 12.92; 4.44, 15.07, 14.49, 13.55, 12.91; 5.45, 14.91, 14.26, 13.49, 13.10; 5.49, 14.85, 14.22, 13.48, 12.91; 5.54, 14.78, 14.02, 13.43, 12.86; 5.58, 14.80, 14.13, 13.33, 12.85; 6.35, 14.96, 14.73, 13.97, 14.01; 6.38, 14.97, 14.29, 13.88, 14.16.

1993 September 16 Daniel W. E. Green